A 13-year bond with a face value of 1500 dollars is redeemable at twice par, and pays annual coupons at a rate of 3.3 percent effective. Find the price to yield an investor 5.9 percent effective.
Answer =  dollars.
Answer #1

coupon payment = 3.3% * 1500 = 49.50

Amount received at maturity = 1500 * 2 = 3000

Present value of bond = 49.50 * (P/A, 5.9%,13) + 3000 * (P/F, 5.9%,13)

= 49.50 * 8.904626 + 3000 * 0.474627

= 1864.66 ~ 1865 (Nearest Dollar)